Sam is a retired teacher who lives in New York City and teaches tennis lessons for extra cash. At a wage of $35 per hour, he is willing to teach 10 hours per week. At $60 per hour, he is willing to teach 19 hours per week.

Using the midpoint method, the elasticity of Sam’s labor supply between the wages of $35 and $60 per hour is approximately _______(.06,.85,1.18,40.28) , which means that Sam’s supply of labor over this wage range is ________(elastic,inelastic) .
Answer #1

Answer:

Given

At $35 / hour Sam teaches for 10 hours

At $60/ hour Sam teaches for 19 hour

Percentage change in supply Q =(19-10)*2/(10+19)=0.62

Percentage change in fees F =(60-35)*2/(35+60)=0.5263

So elasticity of supply =Q/F=0.62/0.5263=1.18

since elasticity of supply is greater than 1 so supply of labor is elastic in this wage range.